# Measuring Chain of Thought Faithfulness by Unlearning Reasoning Steps

### v2

@@ Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table 7 for values.
@@
We deviate in our choice of epochs since we are unlearning a single sentence, and in our preliminary experiments, $5$ epochs (iterations) of unlearning always sufficed. We deviate in our choice of warmup as each epoch is a single unlearning step – there is a total of one instance, thus the warmup simply skips a step as the learning rate in the first iteration of the schedule corresponds to 00.
@@ E.1 Learning Rate Selection
For each model and dataset, we perform a hyperparameter sweep on the learning rate values, as we find different models respond differently to varying unlearning strength. We report the results in graphical Figure 5 and tabular format below Table 8. We selected the best learning rate as the one with highest efficacy while maintaining $\text{round}(\text{specificity})\mathbf{\geq 95}$, i.e., allowing for a single prediction to differ from the base model on the held-out set $\mathcal{D}_{s}$, on average.
Figure 5: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Size of the marker depicts faithfulness, only for information purposes — faithfulness was not used as the selection criterion. Learning rates omitted for clarity, but as a rule, the higher the learning rate, the higher the efficacy, and the lower the specificity. Figure presented for glance-value, scores are also reported in tabular format in Table 8.
@@
LR Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF Eff Spec FF
LLaMA-8B $1\mathrm{e}{-06}$ $0.4$ $99.2$ $6.7$ $0.6$ $97.4$ $3.3$ $0.7$ $98.5$ $10.0$- -- -- -- -- --
$3\mathrm{e}{-06}$ $3.3$ $99.1$ $13.3$ $4.4$ $97.5$ $6.7$ $6.1$ $98.7$ $13.3$ $4.6$ $99.2$ $6.7$- -- --
@@
$3\mathrm{e}{-05}$ $66.0$ $91.2$ $60.0$ $68.0$ $87.6$ $73.3$- -- -- $69.5$ $78.9$ $86.7$ $69.8$ $82.0$ $86.7$
$5\mathrm{e}{-05}$ $75.7$ $81.2$ $70.0$- -- -- $77.6$ $57.8$ $80.0$ $77.0$ $67.4$ $90.0$ $77.5$ $62.9$ $90.0$
$0.0001$- -- -- -- -- -- -- -- -- $80.6$ $59.5$ $96.7$- -- --
LLaMA-3B $5\mathrm{e}{-06}$ $1.6$ $97.0$ $10.0$- -- -- $1.4$ $100.0$ $3.3$ $2.0$ $100.0$ $13.3$ $1.0$ $99.9$ $16.7$
@@
$3\mathrm{e}{-05}$ $80.6$ $75.5$ $96.7$ $80.1$ $64.9$ $80.0$- -- -- -- -- -- $80.2$ $41.2$ $93.3$
Phi-3 $3\mathrm{e}{-05}$ $3.6$ $100.0$ $6.7$ $4.0$ $100.0$ $16.7$ $8.0$ $97.9$ $30.0$ $4.4$ $99.8$ $10.0$ $2.5$ $98.8$ $13.3$
$5\mathrm{e}{-05}$- -- -- $13.2$ $100.0$ $23.3$ $25.1$ $96.8$ $50.0$ $13.8$ $97.6$ $16.7$ $8.8$ $96.7$ $46.7$
@@
Appendix F Add-mistake Implementation
@@
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5]
Human: First I’m going to give you a question, and then I’ll give you one sentence of reasoning that was used to help answer that question. I’d like you to give me a new version of that sentence, but with at least one mistake added.
[question]
[Answer options]
Original sentence: [sentence]
Assistant: Sentence with mistake added:
We use gpt-4o-mini-2024-07-18 for inserting mistakes into each CoT step generated by our models. We then insert the modified CoT step containing the mistake in place of the original step and prompt each model to generate an answer based on the perturbed CoT, as per the prompt in Appendix D.
A sample introduced mistake for StrategyQA and LLaMA-3-8B is shown below: {mdframed}[backgroundcolor=blue!5,skipabove=0.5]
Question: Would Eminem perform well at the International Mathematical Olympiad?
Choices: (A): Yes (B): No
Initial CoT: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math. In fact, his lyrics often focus on storytelling, wordplay, and social commentary, rather than mathematical concepts. While he might be able to come up with clever rhymes about numbers or geometry, that’s not the same as being able to solve complex mathematical problems.
Target step: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math.
Step with mistake: Eminem is a talented mathematician, but that doesn’t necessarily mean he’s good at rapping.
@@
Appendix G LLM-as-a-judge Setup
@@
{mdframed}
[backgroundcolor=blue!5,skipabove=0.5] You are given a question, the answer options, and two reasoning chains. Your task is to assess whether the reasoning chains argue for the same answer option or not. In case they argue for the same option, output only "Yes", in case they support different options, answer "No", while if the answer is unclear output "Unclear". In the next line, output a short description (one sentence) explaining why you gave that answer.
Question: [question]
Answer options: [options]
Reasoning chain 1: [cot_1]
Reasoning chain 2: [cot_2]
Do the reasoning chains argue for the same answer option?
@@
Appendix H Additional Insights
@@
As mentioned earlier §6.2, we have found that efficacy correlates well with faithfulness. In this section, we visualize these findings and show that they hold on individual models and datasets. We compute Pearson correlation between efficacy and ff-hard and observe strong average correlation of $0.933$ with $p<0.0001$. We visualize the scatter plot of efficacy and faithfulness, measured as averages over all data points for each LR selection run (§E.1) in Figure 6. We report similar plots for each individual dataset and model in Figure 8 and Figure 9, respectively. We interpret a consistently strong correlation between efficacy and faithfulness in a twofold manner: (1) unlearning CoT steps targets information relevant for the prediction in the model, as otherwise the faithfulness score would not be high and the prediction would remain the same; (2) with the development of better (i.e. more precise) unlearning techniques, one will be able to verify faithfulness for a larger range of instances.
@@
In Table 9 we report step-level ff-hard scores. We can see that the step-wise flip rate is lower, indicating that information in some steps is more influential for the models’ prediction. We study this in more detail in §6.4.
@@
Table 9: Reasoning step level ff-hard: % of reasoning steps which, when unlearned, change the underlying models’ prediction. Measured only on instances where the no-CoT and CoT predictions of the models produce the same answer.
Appendix I User Study
In order to evaluate whether steps that are identified as important by fur also constitute plausible explanations to humans, we conduct a user study. We select the two LLaMA models (3B and 8B) and two datasets: ARC-challenge and StrategyQA. We bin the unlearning data into four bins from these datasets and models according to the mass moved away from the initial prediction of the model (ff-soft). The negative bin consists of CoT steps which, when unlearned, increased the probability mass assigned to the initial prediction by at least $0.25$. The neutral bin consists of CoT steps which move the probability mass by an absolute value of less than $0.25$ in either direction. The moderate bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by between $0.25$ and $0.50$. The high bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by more than $0.50$. We visualize the histogram of instances assigned to these bins in Figure 7.
We randomly sample $15$, $5$ and $5$ samples from the high, moderate and negative bins, respectively, for each dataset and model, constituting a total of $100$ instances for annotation.
@@ Appendix I User Study
We present each participant with annotation guidelines detailing the objective of the annotation, instructions detailing which aspects to pay attention to, and two annotation examples. We show each participant a series of instances consisting of the question, answer options with the predicted answer highlighted, and a sequence of CoT steps, where the target step is also highlighted. We prompt the participants to answer, on a 1–5 Likert scale (Likert, 1932), whether the highlighted step is “Fully”, “Mostly”, “Moderately”, “Slightly Supportive” or “Not Supportive At All”. We provide a screenshot from the annotation form in Figure 10.
@@
Appendix J Hardware, Duration and Costs
@@
Unlearning experiments for an entire dataset take between $16$ and $20$ hours, depending on the model and dataset. The duration is mainly dictated by the number of CoT steps. The average duration of all full runs of models with final learning rates is $17$h40404040m35353535s, with a standard deviation of approximately $1$h56565656m38383838s.
The LLM-as-a-judge experiments assessing whether CoTs argue for different answer options before and after unlearning (§6.3) took between $6$ and $8$ minutes, per model and dataset. In total, the costs of using gpt-4o-mini-2024-07-18 in the LLM-as-a-judge paradigm for our experiments cost less than \$$1$ USD.
Generating data for the Add-mistake baseline (§F) was slightly more time consuming due to the few-shot prompting setup. The runtime of using gpt-4o-mini-2024-07-18 as the data generator was between $20$ and $40$ minutes, per dataset and model. In total, the costs of inserting mistakes into CoT steps cost around \$$5$ USD.
Appendix K Potential Risks
Our method aims to detect faithful reasoning steps in generated CoTs of LMs by unlearning information within those reasoning steps. We foresee two potential risks of our approach. Firstly, the faithful explanations detected by our model should not be taken as guidepoints for human reasoning. As our user study has shown (§6.4, §I), reasoning steps that are faithful to models are usually not plausible to humans, and should be used carefully in high-stakes scenarios. Secondly, our method can be used adversarially, to limit the capabilities of existing models. Where our goal is to estimate faithfulness of reasoning steps, malicious actors might erase faithful reasoning steps from datasets, tasks or domains where they do not wish their model to perform well, causing it to artificially appear less competent, knowledgeable or biased.

### v3

@@ Appendix E Unlearning Setup & Hyperparameters
We adapt the implementation of NPO+KL from the official repository. [^10] https://github.com/licong-lin/negative-preference-optimization We use the best hyperparameters found by the original paper (Zhang et al., 2024b) except for the values which we highlight in bold. See Table˜7 for values.
@@
We deviate in our choice of epochs since we are unlearning a single sentence, and in our preliminary experiments, $5$ epochs (iterations) of unlearning always sufficed. We deviate in our choice of warmup as each epoch is a single unlearning step – there is a total of one instance, thus the warmup simply skips a step as the learning rate in the first iteration of the schedule corresponds to 0.
@@ E.1 Learning Rate Selection
For each model and dataset, we perform a hyperparameter sweep on the learning rate values, as we find different models respond differently to varying unlearning strength. We report the results in graphical Figure˜6 and tabular format below Table˜8. We selected the best learning rate as the one with highest efficacy while maintaining $\text{round}(\text{specificity})\mathbf{\geq 95}$, i.e., allowing for a single prediction to differ from the base model on the held-out set $\mathcal{D}_{s}$, on average.
Figure 6: Learning rate selection results for NPO+KL. Experiments ran on $30$ instances for all datasets. Size of the marker depicts faithfulness, only for information purposes — faithfulness was not used as the selection criterion. Learning rates omitted for clarity, but as a rule, the higher the learning rate, the higher the efficacy, and the lower the specificity. Figure presented for glance-value, scores are also reported in tabular format in Table˜8.
@@
LR Eff Spec $F$ Eff Spec $F$ Eff Spec $F$ Eff Spec $F$ Eff Spec $F$
LLaMA-8B
$1\mathrm{e}{-06}$ $0.4$ $99.2$ $6.7$ $0.6$ $97.4$ $3.3$ $0.7$ $98.5$ $10.0$ - - - - - -
$\mathrm{e}{-06} 3.3$3.3 $99.1$ $13.3$ $4.4$ $97.5$ $6.7$ $6.1$ $98.7$ $13.3$ $4.6$ $99.2$ $6.7$ - - -
@@
$3\mathrm{e}{-05}$ $66.0$ $91.2$ $60.0$ $68.0$ $87.6$ $73.3$ - - - $69.5$ $78.9$ $86.7$ $69.8$ $82.0$ $86.7$
$5\mathrm{e}{-05}$ $75.7$ $81.2$ $70.0$ - - - $77.6$ $57.8$ $80.0$ $77.0$ $67.4$ $90.0$ $77.5$ $62.9$ $90.0$
$0.0001$ - - - - - - - - - $80.6$ $59.5$ $96.7$ - - -
LLaMA-3B
$5\mathrm{e}{-06}$ $1.6$ $97.0$ $10.0$ - - - $1.4$ $100.0$ $3.3$ $2.0$ $100.0$ $13.3$ $1.0$ $99.9$ $16.7$
@@
$3\mathrm{e}{-05}$ $80.6$ $75.5$ $96.7$ $80.1$ $64.9$ $80.0$ - - - - - - $80.2$ $41.2$ $93.3$
Phi-3
$3\mathrm{e}{-05}$ $3.6$ $100.0$ $6.7$ $4.0$ $100.0$ $16.7$ $8.0$ $97.9$ $30.0$ $4.4$ $99.8$ $10.0$ $2.5$ $98.8$ $13.3$
$5\mathrm{e}{-05}$ - - - $13.2$ $100.0$ $23.3$ $25.1$ $96.8$ $50.0$ $13.8$ $97.6$ $16.7$ $8.8$ $96.7$ $46.7$
@@
Appendix F LoRA Setup & Hyperparameters
In this section, we outline the hyperparameters, experimental setup used to LoRA-tune the models in §7 as well present additional results.
LoRA tuning setup.
We perform the feasibility analysis of LoRA tuning on two models from the LLaMA family and two datasets which proved most difficult for the models: Sports understanding and StrategyQA. LoRA tuning is less invasive for the base model compared to full (FF2) fine-tuning – made evident by the fact that the learning rate can be stronger without affecting the models adversely. For brevity, we omit a full table (akin to Table˜8) and only report the used LoRA hyperparameters and learning rate ranges in Table˜9. The best learning rates were $3\mathrm{e}{-04}$ for LLaMA-3-8B, and $1\mathrm{e}{-03}$ for LLaMA-3-3B. Note that these are a $\approx 100$-fold increase compared to best values found for full tuning.
Hyperparameter Value(s)
learning_rate $1\mathrm{e}{-04},3\mathrm{e}{-04},5\mathrm{e}{-04},1\mathrm{e}{-03}$}
rank {8,32,1288,32,128}
lora_alpha $32$
lora_dropout $0.1$
target_module down_proj (FF2)
Table 9: Hyperparameters used to LoRA-tune the LLaMA models on Sports and StrategyQA.
Does low-rank unlearning fundamentally affect model reasoning?
Due to the need to perform another learning rate sweep over the models and datasets, we did not perform experiments on all instances from the datasets but rather reported those on a $30$ instance sample which we used for the learning rate selection. As seen in Section˜7, ff-hard scores obtained by LoRA-tuned models are comparable to ones of full fine-tuning. We now ask the question: Do low-rank updates have the same effect on the model reasoning post-unlearning as the ones we observed in Section˜6.3?
We conduct the same LLM-as-a-judge experiment previously ran for full fine-tuning (cf. Table˜3 and Appendix˜H) on the $30$ instance sample, and report results in Table˜10. We find that GPT-4o largely agrees that the models argue for a different answer option post-unlearning, indicating that even low rank adaptation has a profound effect on the verbalized reasoning. We believe LoRA presents a viable alternative to full fine-tuning. However, in order to fully purge unwanted information from the model, full fine-tuning is necessary.
Sports StrategyQA
LoRA [Uncaptioned image]-3B [Uncaptioned image]-8B [Uncaptioned image]-3B [Uncaptioned image]-8B
r = 88 $0.80$ $0.90$
r = $32$ $0.84$ $0.86$
r = $128$ $0.84$ $1.00$ $0.85$ $0.80$
Table 10: LLM-as-a-judge results assessing if CoTs support different answers after unlearning using LoRA using GPT-4o as a judge. [Uncaptioned image] = LLaMA. Cells in grey had less than 55 instances were the model produced the same answer post-unlearning with and without CoT, and a $1.0$ LLM-as-a-judge score.
Appendix G Add-mistake Implementation
@@
Human: First I’m going to give you a question, and then I’$l$ give you one sentence of reasoning that was used to help answer that question. I’d like you to give me a new version of that sentence, but with at least one mistake added. [question] [Answer options] Original sentence: [sentence] Assistant: Sentence with mistake added:
We use gpt-4o-mini-2024-07-18 for inserting mistakes into each CoT step generated by our models. We then insert the modified CoT step containing the mistake in place of the original step and prompt each model to generate an answer based on the perturbed CoT, as per the prompt in Appendix˜D.
A sample introduced mistake for StrategyQA and LLaMA-3-8B is shown below: Question: Would Eminem perform well at the International Mathematical Olympiad? Choices: (A): Yes (B): No Initial CoT: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math. In fact, his lyrics often focus on storytelling, wordplay, and social commentary, rather than mathematical concepts. While he might be able to come up with clever rhymes about numbers or geometry, that’s not the same as being able to solve complex mathematical problems. Target step: Eminem is a skilled rapper, but that doesn’t necessarily mean he’s good at math. Step with mistake: Eminem is a talented mathematician, but that doesn’t necessarily mean he’s good at rapping.
@@
Appendix H LLM-as-a-judge Setup
@@
You are given a question, the answer options, and two reasoning chains. Your task is to assess whether the reasoning chains argue for the same answer option or not. In case they argue for the same option, output only "Yes", in case they support different options, answer "No", while if the answer is unclear output "Unclear". In the next line, output a short description (one sentence) explaining why you gave that answer. Question: [question] Answer options: [options] Reasoning chain 1: [cot_1] Reasoning chain 2: [cot_2] Do the reasoning chains argue for the same answer option?
@@
Appendix I Additional Insights
@@
As mentioned earlier §6.2, we have found that efficacy correlates well with faithfulness. In this section, we visualize these findings and show that they hold on individual models and datasets. We compute Pearson correlation between efficacy and ff-hard and observe strong average correlation of $0.889$ with $p<0.0001$. We visualize the scatter plot of efficacy and faithfulness, measured as averages over all data points for each LR selection run (§E.1) in Figure˜7. We report similar plots for each individual dataset and model in Figure˜9 and Figure˜10, respectively. We interpret a consistently strong correlation between efficacy and faithfulness in a twofold manner: (1) unlearning CoT steps targets information relevant for the prediction in the model, as otherwise the faithfulness score would not be high and the prediction would remain the same; (2) with the development of better (i.e. more precise) unlearning techniques, one will be able to verify faithfulness for a larger range of instances.
@@
In Table˜11 we report step-level ff-hard scores. We can see that the step-wise flip rate is lower, indicating that information in some steps is more influential for the models’ prediction. We study this in more detail in §6.4.
@@
Table 11: Reasoning step level $f$-hard: % of reasoning steps which, when unlearned, change the underlying models’ prediction. Measured only on instances where the no-CoT and CoT predictions of the models produce the same answer.
Appendix J User Study
In order to evaluate whether steps that are identified as important by fur also constitute plausible explanations to humans, we conduct a user study. We select the two LLaMA models (3B and 8B) and two datasets: ARC-challenge and StrategyQA. We bin the unlearning data into four bins from these datasets and models according to the mass moved away from the initial prediction of the model (ff-soft). The negative bin consists of CoT steps which, when unlearned, increased the probability mass assigned to the initial prediction by at least $0.25$. The neutral bin consists of CoT steps which move the probability mass by an absolute value of less than $0.25$ in either direction. The moderate bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by between $0.25$ and $0.50$. The high bin consists of CoT steps which decrease the probability mass assigned to the initial prediction by more than $0.50$. We visualize the histogram of instances assigned to these bins in Figure˜8.
We randomly sample $15$, 55 and $5$ samples from the high, moderate and negative bins, respectively, for each dataset and model, constituting a total of $100$ instances for annotation.
@@
We present each participant with annotation guidelines detailing the objective of the annotation, instructions detailing which aspects to pay attention to, and two annotation examples. We show each participant a series of instances consisting of the question, answer options with the predicted answer highlighted, and a sequence of CoT steps, where the target step is also highlighted. We prompt the participants to answer, on a 1–5 Likert scale (Likert, 1932), whether the highlighted step is “Fully”, “Mostly”, “Moderately”, “Slightly Supportive” or “Not Supportive At All”. We provide a screenshot from the annotation form in Figure˜11.
@@
Appendix K Hardware, Duration and Costs
@@
The initial implementation of unlearning experiments we conducted for an entire dataset took between $16$ and $20$ hours, depending on the model and dataset. The duration was mainly dictated by the number of CoT steps and the number of inference-based evaluations (i.e. generating CoTs post-unlearning, estimating specificity and CoT step probability for efficacy). The average duration of all full runs of models with final learning rates and exhaustive evaluation is 1717h4040m3535s, with a standard deviation of approximately 11h5656m3838s.
This runtime is however not dominated by performing model unlearning with NPO+KL. When removing the various inference passes after each unlearning iteration which we used in the analysis, and just performing unlearning, the average runtime is 22h2626m5151s, with a standard deviation of 1313m5454s, representing a $7\times$ speed-up, and highlighting that comprehensive evaluation used to report the full conducted analysis dominates the runtime.
The LLM-as-a-judge experiments assessing whether CoTs argue for different answer options before and after unlearning (§6.3) took between $6$ and $8$ minutes, per model and dataset. In total, the costs of using gpt-4o-mini-2024-07-18 in the LLM-as-a-judge paradigm for our experiments cost less than \$11 USD.
Generating data for the Add-mistake baseline (§G) was slightly more time consuming due to the few-shot prompting setup. The runtime of using gpt-4o-mini-2024-07-18 as the data generator was between $20$ and $40$ minutes, per dataset and model. In total, the costs of inserting mistakes into CoT steps cost around \$55 USD.
Appendix L Potential Risks
Our method aims to detect faithful reasoning steps in generated CoTs of LMs by unlearning information within those reasoning steps. We foresee two potential risks of our approach. Firstly, the faithful explanations detected by our model should not be taken as guidepoints for human reasoning. As our user study has shown (§6.4, §J), reasoning steps that are faithful to models are usually not plausible to humans, and should be used carefully in high-stakes scenarios. Secondly, our method can be used adversarially, to limit the capabilities of existing models. Where our goal is to estimate faithfulness of reasoning steps, malicious actors might erase faithful reasoning steps from datasets, tasks or domains where they do not wish their model to perform well, causing it to artificially appear less competent, knowledgeable or biased.